LHC Report: Good fills but also some technical glitches

During the recent summer weeks the LHC continued running with 1380 bunches, gently pushing up the peak luminosity by reducing the beam sizes and increasing the bunch intensities. This smooth operation was temporarily interrupted by some cryogenics problems and some disturbances to the electrical network.


So far, the LHC has delivered 2.5 fb-1 luminosities to both ATLAS and CMS, giving physicists a steady flow of data to work on. A record peak luminosity (2.4 × 1033 cm-2s-1) was achieved by using smaller-than-nominal beam sizes from the injector chain and increased bunch intensities. The bunch intensities are now up to 1.35 × 1011 protons per bunch, which is well above the nominal value of 1.15 × 1011 protons.

Unfortunately, after a series of good fills, a cold compressor unit stopped working early on Saturday morning, 13 August, causing problems to the cryogenic cooling system at point 8. The system needed about three days of repair and cool-down time in order to fully recover. However, soon after, operation was again disturbed by a site-wide power cut causing the accelerator complex and the associated experiments to stop. It took some time to restore the power manually, and again it was the cryogenics that suffered most from the power outage. Cryogenics specialists worked hard around the clock to re-establish the powering conditions by early Sunday, 21 August.

The first fill after the recovery was used to set the collimation system in order to prepare for further squeezing of the beams at the interaction points in ATLAS and CMS with the aim of increasing the delivered luminosity. These and other tests will continue during the Machine Development period that started on Wednesday 24 August and which will be followed by a week-long technical stop.


by Jan Uythoven for the LHC Team